OpenStudy (anonymous):

Find the total resistance in a parallel dc circuit with two branches rated at 4 and 18, respectively

OpenStudy (anonymous):

Can you give us a drawing? In a parallel circuit the general equation is \[\frac{ 1 }{ R_p} = \frac{ 1 }{ R_1 }+\frac{ 1 }{ R_2 }+...\]

OpenStudy (anonymous):

Where each component has the same voltage